OpenStudy (anonymous):

Complete the inverse of f(x). F(x)= x-6 so, what is f^-1(y)= ?

OpenStudy (anonymous):

y=x-6 swap x and y x=y-6 solve for y y= x+6 f^-1(x) = x+6

OpenStudy (anonymous):

its an inverse, it undoes the other function

OpenStudy (anonymous):

the other one subtracted, this one adds